OpenStudy (anonymous):

find a polynomial function f(x) for the second degree when f(0) =4 , f(-1) = 3 , f(1) =9 .

OpenStudy (anonymous):

f(x) = 2x^2 + 3x +4

OpenStudy (anonymous):

@Vaidehi09 Dont just give the answer. It is against the Code of Conduct. Guide the asker to the answer

OpenStudy (anonymous):

yes please ..

OpenStudy (anonymous):

f(x)=a*x^2+b*x+c we must find a,b and c f(0)=4.....a*0^2+b*0+c=c=4 f(-1)=3....a-b+4=3...a-b=-1 f(1)=9.......a+b+4=9...a+b=5 2a=4and 2b=6 then ; a=2 and b=3 c=4 f(x)=2x^2+3x+4

OpenStudy (anonymous):

here's how: let f(x) = ax^2 + bx + c.........(i) we're given that f(0) = 4. so sub x = 0 in (i) f(0) = 0 + 0 + c 4 = c........(ii) now we know that f(-1) = 3. sub that in (i). f(-1) = a(1) - b + 4.......from (ii) 3 = a - b + 4 b = a+1.........(iii) next, we know f(1) = 9. sub that in (i). f(1) = a + b + 4..........from (ii) 9 = a + a+1 + 4........from (iii) 4 = 2a a = 2. sub this in (iii) u'll get b = 3. sub values of a, b , c in (i) and u'll have ur answer.

OpenStudy (anonymous):

thank you @Vaidehi09

OpenStudy (anonymous):

glad to be of help! :)